OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2x - 14 sin x + 2 = -5

OpenStudy (anonymous):

I have, ...+7=0

OpenStudy (anonymous):

ACtually, I have up to, 7*(sinx-1)(sinx-1)

OpenStudy (anonymous):

So here, you'd want to use the zero-product rule. sin(x) = 1 at what place on a unit circle?

OpenStudy (anonymous):

An example of the zero-product rule, in case: Solve x^2 - 3x + 2 = 0 We first factor: (x - 2) (x - 1) = 0 Then we use the zero product rule to get x - 2 = 0 or x - 1 = 0 Hence x = 2 or x = 1

OpenStudy (anonymous):

Okay, I got sinx=1

OpenStudy (anonymous):

Sweet, so since you still have to solve for x, you find the places on a unit circle where sin is 1. http://www.mathsisfun.com/geometry/unit-circle.html According to the unit circle at the bottom of the page of this link, sin(x)=1 at pi/2

OpenStudy (anonymous):

Oh, okay.

OpenStudy (anonymous):

Thanks.

OpenStudy (anonymous):

Sorry about the late response, but what would the other answers be?

OpenStudy (anonymous):

pi/2, pi/2, and 7=0 is what I got. I'm confused on the 7=0 though.

OpenStudy (anonymous):

Can you please help, anyone?

OpenStudy (anonymous):

What would I do with 7=0?

OpenStudy (mathstudent55):

When you have an equation such as 2(x + 1)(x - 2) = 0, and you use the zero product rule, that only applies to factors that can actually equal zero. 2 = 2 and cannot equal zero, so don't do anything with it. x + 1 = 0 or x - 2 = 0 will give you solutions x = -1 or x = 2. Actually, there is a mathematical thing you can do at this stage: 2(x + 1)(x - 2) = 0 Before applying the zero product rule, divide both sides by 2: (x + 1) (x - 2) = 0 x + 1 = 0 or x - 2 = 0 x = -1 or x = 2 Your case is 7 sin^2x - 14 sin x + 2 = -5 7 sin^2x - 14 sin x + 7 = 0 7(sin^2x - 2 sin x + 1) = 0 7(sin x - 1)(sin x - 1) = 0 Now divide both sides by 7 sin x - 1 = 0 or sin x - 1 = 0 sin x = 1 or sin x = 1 You get x = pi/2 twice, and that's it.